The first attached figure compares acoustic and seismic noise of the NEB experimental hall and tower platform (Guralp sensor) now (purple, Nov 9 2021) and in O3 (blue, Feb 8 2020). The noise reduction in the band few Hz to about 100 Hz is evident in all sensors. Note also that the actual noise level it is essentially everywhere below the O3 noise.
The noise floor significantly reduced, as well as several spectral lines. One for all, the 33 Hz line which was a scattered light offender during O3.
A careful understanding of single contributions is in progress, yet we can already tell that two major mitigation steps have been:
- insertion of ducts soft sleeves, on April 27 2021 - elog 51545: the noise reduction impact on the same sensors is shown in Figure 2,
- slow down of the AHU (from 50 Hz to 40 Hz) on June 8 2021: shown in Figure 3.